# New Lower Bounds for Stochastic Non-Convex Optimization through Divergence Composition

### v1

@@ Appendix E Technical Results
Appendix F Canonical Model and Divergence Composition
@@
We consider an optimization algorithm $\mathcal{A}$ specified by a sequence $\bigl{(}K_{t}\bigr{)}_{1\leq t\leq T}$ of probability kernels $K_{t}:\Omega_{t-1}\to\mathbb{R}^{d_{0}}$, which model the choice of the query point $\mathbf{x}_{t}$ at time $t$. At the first round, $\mathbf{x}_{1}$ is chosen according to a distribution defined by $K_{1}$.
For each $i\in\{0,\dots,m\}$, let $H_{i,t}:\Omega_{t-1}\times\mathbb{R}^{d_{0}}\times\mathbb{R}^{d_{0}}\;\to\;\mathbb{R}^{d_{0}}$ be the probability kernel that models the feedback observed by the learner. For any measurable set $B\subseteq\mathbb{R}^{d_{0}}$, we take
$H_{i,t}(\bm{y}_{t-1},\bm{x}_{t},B)=\left(\frac{d_{0}}{2\pi\sigma^{2}}\right)^{d_{0}/2}\int_{\bm{x}\in B}\exp\left(-\frac{d_{0}}{2\sigma^{2}}\left\|{\bm{x}-\nabla f_{i}(\bm{x}_{t})}\right\|^{2}\right)\mathop{}\!\mathrm{d}\bm{x}~{}.$
@@
$\text{KL}\left(\mathbb{P}_{0}^{T},\mathbb{P}_{i}^{T}\right)$ $\leq\frac{d_{0}}{2\sigma^{2}}\mathbb{E}_{0}[N_{S}]\sup_{\bm{x}\in S}\ \left\|{\nabla f_{0}(\bm{x})-\nabla f_{i}(\bm{x})}\right\|^{2}$
$\qquad+\frac{d_{0}}{2\sigma^{2}}\mathbb{E}_{0}[T-N_{S}]\sup_{\bm{x}\notin S}\ \left\|{\nabla f_{0}(\bm{x})-\nabla f_{i}(\bm{x})}\right\|^{2}~{}.$
@@
$\text{KL}(\mathbb{P}_{0,t},\mathbb{P}_{1,t})$ $=\text{KL}\left(H_{0,t}\circ K_{t}\circ\mathbb{P}_{0,t-1},H_{i,t}\circ K_{t}\circ\mathbb{P}_{i,t-1}\right)$
$=\text{KL}(K_{t}\circ\mathbb{P}_{0,t-1},K_{t}\circ\mathbb{P}_{1,t-1})$
$\qquad+\int_{\Omega_{t-1}\times\mathbb{R}^{d_{0}}}\text{KL}\left(H_{0,t}(\bm{y}_{t-1},\bm{x}_{t},\cdot),H_{i,t}(\bm{y}_{t-1},\bm{x}_{t},\cdot)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\bm{y}_{t-1},\bm{x}_{t})$
@@
$\quad+\int_{\Omega_{t-1}\times\mathbb{R}^{d_{0}}}\text{KL}\left(H_{0,t}(\bm{y}_{t-1},\bm{x}_{t},\cdot),H_{i,t}(\bm{y}_{t-1},\bm{x}_{t},\cdot)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\bm{x}_{t})$
@@
$\quad+\int_{\mathbb{R}^{d_{0}}}\text{KL}\left(\mathcal{N}_{d_{0}}\left(\nabla f_{0}(\bm{x}_{t}),\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\right),\mathcal{N}_{d_{0}}\left(\nabla f_{i}(\bm{x}_{t}),\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\right)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\bm{x}_{t})$
@@
where we used in the third line the fact that given a fixed history input $\bm{y}_{t-1}$, the selected point $\bm{x}_{t}$ by $\mathcal{A}$ have the same distribution regardless of the objective function, and the penultimate inequality comes from the fact that the density of the kernel $H_{i,t-1}$ depends only on the last coordinate $\bm{x}_{t}$, and is exactly that of a Gaussian variable.
@@
$\leq\mathbb{P}_{0}\left(\bm{X}_{t}\in S\right)\sup_{\bm{x}\in S}\left\|{\nabla f_{0}(\bm{x})-\nabla f_{i}(\bm{x})}\right\|^{2}$
$\qquad+\mathbb{P}_{0}\left(\bm{X}_{t}\notin S\right)\sup_{\bm{x}\notin S}\left\|{\nabla f_{0}(\bm{x})-\nabla f_{i}(\bm{x})}\right\|^{2}~{}.$ (21)

### v2

@@ Appendix E Technical Results
Appendix F Canonical Model and Divergence Decomposition
@@
We consider an optimization algorithm $\mathcal{A}$ specified by a sequence $\bigl{(}K_{t}\bigr{)}_{1\leq t\leq T}$ of probability kernels $K_{t}:\Omega_{t-1}\times\mathcal{B}(\mathbb{R}^{d_{0}})\to[0,1]$, which model the choice of the query point $\mathbf{x}_{t}$ at time $t$. More precisely, in round $t$ given the history $\mathbf{y}_{t-1}\in\Omega_{t-1}$, the distribution of $\mathbf{x}_{t}$ is given using the probability measure $A\mapsto K_{t}(\mathbf{y}_{t-1},A)$. At the first round, $\mathbf{x}_{1}$ is chosen according to a distribution defined by $K_{1}$.
For each $i\in\{0,\dots,m\}$, let $H_{i,t}:\Omega_{t-1}\times\mathbb{R}^{d_{0}}\times\mathcal{B}\left(\mathbb{R}^{d_{0}}\right)\;\to\;[0,1]$ be the probability kernel that models the feedback observed by the learner. For any measurable set $B\in\mathcal{B}\left(\mathbb{R}^{d_{0}}\right)$, we take
$H_{i,t}(\bm{y}_{t-1},\mathbf{x}_{t},B)=\left(\frac{d_{0}}{2\pi\sigma^{2}}\right)^{d_{0}/2}\int_{\mathbf{x}\in B}\exp\left(-\frac{d_{0}}{2\sigma^{2}}\left\|{\mathbf{x}-\nabla f_{i}(\mathbf{x})}\right\|^{2}\right)\mathop{}\!\mathrm{d}\mathbf{x}~{}.$
@@
$\text{KL}\left(\mathbb{P}_{0}^{T},\mathbb{P}_{i}^{T}\right)$ $\leq\frac{d_{0}}{2\sigma^{2}}\mathbb{E}_{0}[N_{S}]\sup_{\mathbf{x}\in S}\ \left\|{\nabla f_{0}(\mathbf{x})-\nabla f_{i}(\mathbf{x})}\right\|^{2}$
$\qquad+\frac{d_{0}}{2\sigma^{2}}\mathbb{E}_{0}[T-N_{S}]\sup_{\mathbf{x}\notin S}\ \left\|{\nabla f_{0}(\mathbf{x})-\nabla f_{i}(\mathbf{x})}\right\|^{2}~{}.$
@@
$\text{KL}(\mathbb{P}_{0,t},\mathbb{P}_{i,t})$ $=\text{KL}\left(H_{0,t}\circ K_{t}\circ\mathbb{P}_{0,t-1},H_{i,t}\circ K_{t}\circ\mathbb{P}_{i,t-1}\right)$
$=\text{KL}(K_{t}\circ\mathbb{P}_{0,t-1},K_{t}\circ\mathbb{P}_{i,t-1})$
$\qquad+\int_{\Omega_{t-1}\times\mathbb{R}^{d_{0}}}\text{KL}\left(H_{0,t}(\mathbf{y}_{t-1},\mathbf{x}_{t},\cdot),H_{i,t}(\mathbf{y}_{t-1},\mathbf{x}_{t},\cdot)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\mathbf{y}_{t-1},\mathbf{x}_{t})$
@@
$\quad+\int_{\Omega_{t-1}\times\mathbb{R}^{d_{0}}}\text{KL}\left(H_{0,t}(\mathbf{y}_{t-1},\mathbf{x}_{t},\cdot),H_{i,t}(\mathbf{y}_{t-1},\mathbf{x}_{t},\cdot)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\mathbf{y}_{t-1},\mathbf{x}_{t})$
@@
$\quad+\int_{\Omega_{t-1}\times\mathbb{R}^{d_{0}}}\text{KL}\left(\mathcal{N}_{d_{0}}\left(\nabla f_{0}(\mathbf{x}_{t}),\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\right),\mathcal{N}_{d_{0}}\left(\nabla f_{i}(\mathbf{x}_{t}),\frac{\sigma^{2}}{d_{0}}\bm{I}_{d_{0}}\right)\right)\bm{\mathrm{d}}K_{t}\circ\mathbb{P}_{0,t-1}(\mathbf{y}_{t-1},\mathbf{x}_{t})$
@@
where we used in the third line the fact that the algorithm $\mathcal{A}$ is fixed and the learner internal randomization is independent of the objective function, therefore given a fixed history input $\mathbf{y}_{t-1}$ we have $K(\cdot\mid\mathbf{y}_{t-1})$ is identical under the two hypotheses. The penultimate inequality comes from the fact that the density of the kernel $H_{i,t-1}$ depends only on the last coordinate $\bm{x}_{t}$, and is exactly that of a Gaussian variable.
@@
$\leq\mathbb{P}_{0}\left(\bm{X}_{t}\in S\right)\sup_{\mathbf{x}\in S}\left\|{\nabla f_{0}(\mathbf{x})-\nabla f_{i}(\mathbf{x})}\right\|^{2}$
$\qquad+\mathbb{P}_{0}\left(\bm{X}_{t}\notin S\right)\sup_{\mathbf{x}\notin S}\left\|{\nabla f_{0}(\mathbf{x})-\nabla f_{i}(\mathbf{x})}\right\|^{2}~{}.$ (21)